OpenStudy (anonymous):

please help i suck at math!! ill give a medal & become a fan too! 6C3 A. 120 B. 720 C. 20 D. 240

OpenStudy (anonymous):

@LOUISndGUCCI can u help? please

OpenStudy (anonymous):

whats the question/

OpenStudy (anonymous):

@maddiegirl

OpenStudy (anonymous):

i have to solve 6C3

OpenStudy (anonymous):

its c So, 6C3 is: 6! divided by (3! x (6-3)!) that is, 6! / (3! x 3!) that is, (6 x 5 x 4) / (3 x 2 x 1) = 20 @maddiegirl

OpenStudy (anonymous):

okay thanks! could u help with some others?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@maddiegirl just ask away

OpenStudy (anonymous):

okay Solve: 6P4 A. 240 B. 720 C. 20 D. 360

OpenStudy (anonymous):

6P4 = 6! / (6-4)! = 6! / 2! = 6 x 5 x 4 x 3 x 2 x 1 / 2 x 1 = 360

OpenStudy (anonymous):

thanks tag me whenever needed do you need more help?

OpenStudy (anonymous):

THANK U! do u got time for a few more?

OpenStudy (anonymous):

& okay i will & yes

OpenStudy (anonymous):

i got time u got those right correct?

OpenStudy (anonymous):

okay & yes i think so

OpenStudy (anonymous):

How many ways can you choose 3 dishes from a menu of 8? A. 24(i picked this) B. 336 C. 144 D. 56 How many different ways could you arrange 4 books on a shelf? A. 4 B. 24 C. 14 D. 8 How many different ways can you choose 3 frozen yogurt toppings from 20? A. 6,840 B. 1,140 C. 60(i picked this) D. 2,260 Your class is electing a president and vice president. How many ways can a president and vice president be selected from 25? A. 300 B. 150 C. 600 D. 50(i picked this)

OpenStudy (anonymous):

@LOUISndGUCCI

OpenStudy (anonymous):

@taylor12344 could u help?